### v1

@@ Abstract
The aging population poses numerous challenges to healthcare, including the increase in chronic wounds in the elderly. The current approach to wound assessment by therapists based on photographic documentation is subjective, highlighting the need for computer-aided wound recognition from smartphone photos. This offers objective and convenient therapy monitoring, while being accessible to patients from their home at any time. However, despite research in mobile image segmentation, there is a lack of focus on mobile wound segmentation. To address this gap, we conduct initial research on three lightweight architectures to investigate their suitability for smartphone-based wound segmentation. Using public datasets and UNet as a baseline, our results are promising, with both ENet and TopFormer, as well as the larger UNeXt variant, showing comparable performance to UNet. Furthermore, we deploy the models into a smartphone app for visual assessment of live segmentation, where results demonstrate the effectiveness of TopFormer in distinguishing wounds from wound-coloured objects. While our study highlights the potential of transformer models for mobile wound segmentation, future work should aim to further improve the mask contours{extended} [^1] This is an extended version of our KI 2024 paper, available at the Springer website..
@@ 3.1 Dataset
We perform experiments on a combination of two prominent datasets: the Foot Ulcer Segmentation Challenge 2021 (FuSeg) [34] and the Diabetic Foot Ulcer Challenge 2022 (DFUC 2022) [13, 36] datasets. FuSeg encompasses 1,210 foot ulcer images, each with a resolution of 512 × 512 pixels, sourced from 889 patients’ medical visits spanning from October 2019 to April 2021. Meanwhile, the DFUC 2022 dataset comprises 4,000 foot ulcer images with a resolution of 640 × 480 pixels, captured approximately 30 to 40 cm away from the ulcer. Our custom dataset, Combined Foot Ulcers (CFU), integrates FuSeg and DFUC 2022, comprising images with mixed resolutions of either 512 × 512 or 640 × 480 pixels. Due to the privacy of test set annotations (200 out of 1,210 for FuSeg and 2,000 out of 4,000 for DFUC 2022), only publicly accessible training and validation images are utilized for CFU. To prevent data leakage, duplicate images within FuSeg and DFUC 2022 were identified and removed. A similarity analysis, employing the perceptual hashing algorithm [18], was conducted. Image sets with identical raw bytes or perceptual hash values exhibiting a Hamming distance of 11 or less were deemed duplicates. {short}For illustrative purposes, a number of examples are provided in the appendix of the extended version of this paper. {extended}For illustrative purposes, a number of examples are provided in Figure 2. Eliminating duplicate image-annotation pairs, we obtained a final dataset with 2,887 samples after discarding 123 pairs. Next, these samples were randomly partitioned into training (60%), validation (20%), and test (20%) set.
@@ 3.3 Implementation Details
We implemented this approach using Python 3.10.9 and PyTorch 2.0.1 on a server with two partitioned Nvidia A100 GPUs, reserving a large-sized instance with  40 GB VRAM for our experiments. All models were trained end-to-end using the AdamW optimizer with an initial learning rate of 0.0001, a batch size of four, and the binary cross entropy (BCE) loss. We used a ReduceLROnPlateau scheduler conditioned on the Dice score, with a minimum learning rate of 0.000001. All models were trained for 200 epochs from scratch and for at least 130 epochs in the pre-training setting. For both, the best checkpoints, according to the mean IoU on the validation set, were used for the ensuing test set evaluation. All images were resized to 512 × 512 pixels and normalized during training. We applied thresholding to annotated masks to eliminate out-of-class pixels and used the following data augmentations for each training set sample: Gaussian blur with a kernel size of 25 and a randomly sampled standard deviation from a uniform distribution of [0.001, 2.0], random affine transforms (incl. translations by up to 12.5%, rotations by up to 180 degrees, scaling between 50% and 150% of the original size, and shear of up to 22.5 degrees), color jitter, as well as random horizontal and vertical flips, each with a 50% probability. {short}Further details regarding the augmentations are appended in the extended version in the form of an example configuration file. {extended}Details about the augmentations can be found in the example configuration file of Listing 0.A.
@@ 4.1 Results
We compare the selected models in Table 2, where we report the micro-averaged metric scores of each variant for both training modalities. Our findings indicate that, without pretraining, UNeXt-B exhibits the most favorable segmentation performance, balancing parameter count and segmentation quality. Notably, all lightweight models, except UNeXt-S, achieve results within a similar size range as the larger UNet, with the base variants of both UNeXt and Topformer, and ENet even slightly surpassing UNet. Pretraining considerably enhances performance across all models in terms of IoU and DSC metrics. Specifically, all lightweight networks demonstrate comparable performance, with IoU ranging from 69% to 72% and DSC from 82% to 84%, except for UNeXt-S, which performs worse.

### v3

@@ Abstract
The aging population poses numerous challenges to healthcare, including the increase in chronic wounds in the elderly. The current approach to wound assessment by therapists based on photographic documentation is subjective, highlighting the need for computer-aided wound recognition from smartphone photos. This offers objective and convenient therapy monitoring, while being accessible to patients from their home at any time. However, despite research in mobile image segmentation, there is a lack of focus on mobile wound segmentation. To address this gap, we conduct initial research on three lightweight architectures to investigate their suitability for smartphone-based wound segmentation. Using public datasets and UNet as a baseline, our results are promising, with both ENet and TopFormer, as well as the larger UNeXt variant, showing comparable performance to UNet. Furthermore, we deploy the models into a smartphone app for visual assessment of live segmentation, where results demonstrate the effectiveness of TopFormer in distinguishing wounds from wound-coloured objects. While our study highlights the potential of transformer models for mobile wound segmentation, future work should aim to further improve the mask contours [^1] This is an extended version of our KI 2024 paper, which is available at Springer..
@@ 3.1 Dataset
We perform experiments on a combination of two prominent datasets: the Foot Ulcer Segmentation Challenge 2021 (FuSeg) [34] and the Diabetic Foot Ulcer Challenge 2022 (DFUC 2022) [13, 36] datasets. FuSeg encompasses 1,210 foot ulcer images, each with a resolution of 512 × 512 pixels, sourced from 889 patients’ medical visits spanning from October 2019 to April 2021. Meanwhile, the DFUC 2022 dataset comprises 4,000 foot ulcer images with a resolution of 640 × 480 pixels, captured approximately 30 to 40 cm away from the ulcer. Our custom dataset, Combined Foot Ulcers (CFU), integrates FuSeg and DFUC 2022, comprising images with mixed resolutions of either 512 × 512 or 640 × 480 pixels. Due to the privacy of test set annotations (200 out of 1,210 for FuSeg and 2,000 out of 4,000 for DFUC 2022), only publicly accessible training and validation images are utilized for CFU. To prevent data leakage, duplicate images within FuSeg and DFUC 2022 were identified and removed. A similarity analysis, employing the perceptual hashing algorithm [18], was conducted. Image sets with identical raw bytes or perceptual hash values exhibiting a Hamming distance of 11 or less were deemed duplicates. For illustrative purposes, a number of examples are provided in Figure 2. Eliminating duplicate image-annotation pairs, we obtained a final dataset with 2,887 samples after discarding 123 pairs. Next, these samples were randomly partitioned into training (60%), validation (20%), and test (20%) set.
@@ 3.3 Implementation Details
We implemented this approach using Python 3.10.9 and PyTorch 2.0.1 on a server with two partitioned Nvidia A100 GPUs, reserving a large-sized instance with  40 GB VRAM for our experiments. All models were trained end-to-end using the AdamW optimizer with an initial learning rate of 0.0001, a batch size of four, and the binary cross entropy (BCE) loss. We used a ReduceLROnPlateau scheduler conditioned on the Dice score, with a minimum learning rate of 0.000001. All models were trained for 200 epochs from scratch and for at least 130 epochs in the pre-training setting. For both, the best checkpoints, according to the mean IoU on the validation set, were used for the ensuing test set evaluation. All images were resized to 512 × 512 pixels and normalized during training. We applied thresholding to annotated masks to eliminate out-of-class pixels and used the following data augmentations for each training set sample: Gaussian blur with a kernel size of 25 and a randomly sampled standard deviation from a uniform distribution of [0.001, 2.0], random affine transforms (incl. translations by up to 12.5%, rotations by up to 180 degrees, scaling between 50% and 150% of the original size, and shear of up to 22.5 degrees), color jitter, as well as random horizontal and vertical flips, each with a 50% probability. Details about the augmentations can be found in the example configuration file of Listing 0.A.
@@ 4.1 Results
We compare the selected models in Table 2, where we report the micro-averaged metric scores of each variant for both training modalities. Our findings reveal that, without pretraining, UNeXt-B demonstrates the most favorable segmentation performance in terms of DSC and IoU scores, effectively balancing parameter count and segmentation quality. Notably, all lightweight models, except UNeXt-S, achieve results within a similar size range as the larger-sized UNet, with the base variants of UNeXt and TopFormer, and ENet even slightly surpassing UNet. Pretraining considerably enhances performance across all models regarding IoU and DSC metrics. Specifically, all lightweight networks show comparable performance, with IoU scores ranging from 69% to 72% and DSC values from 82% to 84%, except for UNeXt-S, which performs worse.